# HugAgent: Evaluating LLMs in Simulating Individual-Level Human Reasoning on Open-Ended Tasks

### v2

@@ 8   Realize $(e_{j},b_{j-1},b_{j})$ to text; append to transcript;
  • Task 1: Belief-State Inference. Given a participant’s responses and contextual cues, models must infer the person’s stance and factor-level attribution. Example questions include: “Does the respondent view low-income housing as a positive or negative effect on property values?”
@@
Figure 7: Overview of the HugAgent benchmark structure. Inputs (demographics, questionnaires, and transcripts) are mapped to outputs, including belief-state inference (Task 1) and belief dynamics update (Task 2).
@@
Let $y_{i}$ denote the ground-truth response for instance $i$, $\hat{y}_{i}$ the model prediction, and $N$ the total number of instances. For belief dynamic update tasks, let $y_{i}^{\text{prev}}$ denote the participant’s pre-intervention score.
@@
We define directional accuracy as a two-stage weighted metric: (1) detecting whether a belief change occurred, and (2) if so, correctly predicting the direction of change (increase, decrease, or no change). To better reflect the importance of directional reasoning, a higher weight is assigned to the second stage.
$\text{DirAcc}=\lambda\cdot\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}\!\left[\,\Delta\hat{y}_{i}=0\wedge\Delta y_{i}=0\,\right]+(1-\lambda)\cdot\frac{1}{|\mathcal{C}|}\sum_{i\in\mathcal{C}}\mathbf{1}\!\left[\,\operatorname{sgn}(\Delta\hat{y}_{i})=\operatorname{sgn}(\Delta y_{i})\,\right],$ (4)
@@
This weighting reflects the intuition that correctly inferring the direction of belief change is more informative than merely detecting whether a change occurred. While the first stage (change detection) captures a coarse perceptual judgment that can often be guessed in noisy or stable settings, the second stage(directional correctness)reveals whether the model truly understands and reasons about belief dynamics. Hence, emphasizing the latter better reflects a model’s fidelity to human-like reasoning processes and its capacity to simulate belief evolution.
Unified Aggregate Score.
To provide a single comprehensive measure of model performance across both static and dynamic belief tasks, we define a unified score that integrates the belief state inference (BSI) and belief dynamic update (BDU) components into a normalized value within $[0,1]$.
Specifically, $S_{\text{BSI}}$ denotes the normalized accuracy score for static belief state inference, while $S_{\text{BDU}}$ aggregates multiple metrics from the belief dynamic update task, including tolerance accuracy, normalized MAE, and directional reasoning accuracy.
The unified score is computed as:
$\text{UnifiedScore}=\tfrac{1}{2}S_{\text{BSI}}+\tfrac{1}{2}\Big[\tfrac{1}{2}\big(\tfrac{1}{2}S_{\text{BDU-mae-norm}}+\tfrac{1}{2}S_{\text{BDU-acc}}\big)+\tfrac{1}{2}S_{\text{Directional-acc}}\Big].$ (5)
@@
where $\text{MAE}_{\max}$ denotes the task-specific upper bound of allowable error.
The unified score assigns equal weights to the state component ($S_{\text{BSI}}$) and the update component ($S_{\text{BDU}}$). Within the update branch, tolerance accuracy and MAE are equally weighted (0.5 each) to ensure a balanced consideration of robustness and precision. The directional reasoning component ($S_{\text{Directional-acc}}$) is assigned an equal weight (0.5) relative to the combined MAE and accuracy branch, reflecting its comparable importance in capturing belief-updating dynamics.
@@
Unless otherwise specified, all quantitative analyses and unified aggregate score (UAS) computations are performed on the human track, which serves as the primary evaluation benchmark due to its ecological validity and authentic reasoning diversity.
@@
In practice, model predictions for both tasks—belief-state inference (BSI) and belief dynamic update (BDU)—are first computed independently. All metrics (accuracy, MAE, and directional accuracy) are averaged across the three domains (healthcare, surveillance, zoning) before aggregation into the unified score defined in Equation 5. Reported results in Section 4 and subsequent findings are therefore based on the human track unless explicitly noted otherwise.

### v3

@@ 8   Realize $(e_{j},b_{j-1},b_{j})$ to text; append to transcript;
  • Task 1: Belief State Inference. Given a participant’s responses and contextual cues, models must infer the person’s stance and factor-level attribution. Example questions include: “Does the respondent view low-income housing as a positive or negative effect on property values?”
@@
Figure 7: Overview of the HugAgent benchmark structure. Inputs (demographics, questionnaires, and transcripts) are mapped to outputs, including belief state inference (Task 1) and belief dynamics update (Task 2).
@@
Let $y_{i}$ denote the ground-truth response for instance $i$, $\hat{y}_{i}$ the model prediction, and $N$ the total number of instances. For belief dynamics update tasks, let $y_{i}^{\text{prev}}$ denote the participant’s pre-intervention score.
@@
We define directional accuracy as a two-stage weighted metric: (1) Change Detection: detecting whether a belief change occurred, and if so, (2) Directional Inference: correctly predicting the direction of change (increase, decrease, or no change). To better reflect the importance of directional reasoning, a higher weight is assigned to the second stage.
$\text{DirAcc}={}$ $\lambda\cdot\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}\!\left[(\Delta y_{i}=0\wedge\Delta\hat{y}_{i}=0)\ \vee\ (\Delta y_{i}\neq 0\wedge\Delta\hat{y}_{i}\neq 0)\right]$ (4)
$+(1-\lambda)\cdot\frac{1}{|\mathcal{C}|}\sum_{i\in\mathcal{C}}\mathbf{1}\!\left[\operatorname{sgn}(\Delta y_{i})=\operatorname{sgn}(\Delta\hat{y}_{i})\right],$
@@
This weighting reflects the intuition that correctly inferring the direction of belief change is more informative than merely detecting whether a change occurred. While the first stage (change detection) captures a coarse perceptual judgment that can often be guessed in noisy or stable settings, the second stage(directional inference)reveals whether the model truly understands and reasons about belief dynamics. Hence, emphasizing the latter better reflects a model’s fidelity to human-like reasoning processes and its capacity to simulate belief evolution.
Average-to-Individual (ATI) Score.
To provide a single comprehensive measure of model performance across both static and dynamic belief tasks, we define a unified score, Average-to-Individual (ATI) score, that integrates the Belief State Inference (BSI) and Belief Dynamics Update (BDU) components into a normalized value within $[0,1]$. Specifically, $S_{\text{BSI}}$ denotes the normalized accuracy score for static belief state inference, while $S_{\text{BDU}}$ aggregates multiple metrics from the belief dynamics update task, including tolerance accuracy, normalized MAE, and directional reasoning accuracy. The unscaled ATI score is computed as:
$\text{ATI}_{\text{unscaled}}=\tfrac{1}{2}S_{\text{BSI}}+\tfrac{1}{2}\Big[\tfrac{1}{2}\big(\tfrac{1}{2}S_{\text{BDU-mae-norm}}+\tfrac{1}{2}S_{\text{BDU-acc}}\big)+\tfrac{1}{2}S_{\text{Directional-acc}}\Big].$ (5)
@@
where $\text{MAE}_{\max}$ denotes the task-specific upper bound of allowable error. The unified score assigns equal weights to the state component ($S_{\text{BSI}}$) and the update component ($S_{\text{BDU}}$). Within the update branch, tolerance accuracy and MAE are equally weighted (0.5 each) to ensure a balanced consideration of robustness and precision. The directional reasoning component ($S_{\text{Directional-acc}}$) is assigned an equal weight (0.5) relative to the combined MAE and accuracy branch, reflecting its comparable importance in capturing belief-updating dynamics.
To facilitate interpretation relative to human performance and baseline behavior, we linearly rescale $\text{ATI}_{\text{unscaled}}$ to a 0–100 scale:
$\text{ATI}=\frac{\text{ATI}_{\text{unscaled}}-\text{ATI}_{\text{random}}}{\text{ATI}_{\text{human}}-\text{ATI}_{\text{random}}}\times 100,$ (7)
where $\text{ATI}_{\text{random}}$ and $\text{ATI}_{\text{human}}$ denote the unscaled ATI scores of the random guess baseline and human upper bound, respectively. Under this rescaling, a score of 0 indicates random-level performance, while 100 represents human-level performance.
@@
Unless otherwise specified, all quantitative analyses and unified average to individual (ATI) score computations are performed on the human track, which serves as the primary evaluation benchmark due to its ecological validity and authentic reasoning diversity.
@@
In practice, model predictions for both tasks—belief state inference (BSI) and belief dynamics update (BDU)—are first computed independently. All metrics (accuracy, MAE, and directional accuracy) are averaged across the three domains (healthcare, surveillance, zoning) before aggregation into the unified score defined in Equation 5. Reported results in Section 4 and subsequent findings are therefore based on the human track unless explicitly noted otherwise.